OpenStudy (anonymous):

differentiate y^3+xy+2=0 (find dy/dx) answer is -2/7 (i think) but not sure how to get there, i got -1/3 can someone differentiate and show the work quick?

OpenStudy (anonymous):

Is this the chapter on implicit differentiation?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

\[3y^2y'+ xy'+y=0 \]

OpenStudy (anonymous):

brb sec, i have a question being answered.

OpenStudy (anonymous):

i got f'(x)= 3y^2(dy/dx)+y+x(dy/dx) 0 = -y= dy/dx (3y^2+x) then divided by (3y^2+x) so got dy/dx = -y/3y^2+x is that right?

OpenStudy (anonymous):

thats what I just got.

OpenStudy (anonymous):

but then the next question is to find the tan line approx to the curve at the point (-5,2)

OpenStudy (anonymous):

plug in the values

OpenStudy (anonymous):

and i can't do that if i dont have an m value (which i think is found in the last question)

OpenStudy (anonymous):

m is your y'

OpenStudy (anonymous):

(y-2) = y'(x+5)

OpenStudy (anonymous):

ohh got it. it wasn't clicking tht the y' was dy/dx

OpenStudy (anonymous):

thanks